# C.V. Raman Avenue, Bengaluru 560012, India

### 18 Dec 2023 v1

@@ 3.3 Mixed field XXZ Hamiltonian
The XXZ Hamiltonian commutes with total spin operator $M$ in the $z$-direction and is also invariant under reflection with respect to the edge of the chain i.e. under parity operator $P$. In previous studies Santos_2004 ; PhysRevE.84.016206 ; PhysRevB.80.125118 ; PhysRevB.98.235128 ; PhysRevB.102.075127 ; PhysRevX.10.041017 ; gubin_quantum_2012 ; rabinovici_krylov_2022 , it was shown that the integrability can be broken by addition local term
@@
$\mathcal{H}^{(\alpha)}_{XXZ}=\frac{1}{4}\sum_{i=1}^{N-1}J^{\alpha}_{ij}\,(\sigma_{i}^{x}\sigma_{i+1}^{x}+\sigma_{i}^{y}\sigma_{i+1}^{y})+\frac{\tilde{J}^{\alpha}_{ij}}{4}\,\sigma_{i}^{z}\sigma_{i+1}^{z}$ (18)
@@
We study the initial Lanczos sequence for $L=12$ with varying exponent $\alpha$ for zero and non-zero value of $\epsilon_{d}$ (See Fig. 7). The initial operator $\mathcal{O}$ is chosen such that it respects the two symmetries of the model, i.e., parity and total spin in $z$-direction
@@ 4 Conclusions
4.1 Discussion of results
Lanczos sequence
The saturation and initial growth of Lanczos coefficients is greater in non-local TFIM Hamiltonian relative to the local case, which reflects the faster information scrambling in non-local systems. The local chaotic dynamics, along with long-range non-local interactions, are sufficient to give rise to fast scrambling. This is because the Lanczos sequence for chaotic Hamiltonian with local interactions saturates at a comparatively smaller value than its non-local chaotic counterpart. The large saturation value of Lanczos coefficients are the result of non-local interactions in the system, but they can equivalently be said to be the result of fast scrambling as well. The fluctuation in the non-local cases is also suppressed, which is a result of integrability breaking due to non-local interaction. As we see later, this results in a higher saturation value of Krylov complexity.
Krylov Complexity
Apart from common features of Krylov complexity, such as the transition from exponential growth to linear and the late time saturation, we find that the initial time growth rate is greater in non-local cases, showing the fast scrambling in the system. The higher growth rate in Krylov complexity for the non-local cases is a result of the initial faster growth of Lanczos coefficients in non-local cases, thus indicating faster scrambling due to the non-local interaction. Furthermore, since non-locality results in integrability breaking, we also find a large saturation value in Krylov complexity satisfying the bound in terms of the Krylov dimension. It is important to note that although the non-local interactions with the integrable local part of Hamiltonian show level statistics that of a chaotic model, it does not attain K-complexity saturation value similar to that of a chaotic case. In terms of associated Lanczos coefficients, the non-local integrable case has larger fluctuations as compared to chaotic cases.
@@ 4.1 Discussion of results
4.2 Outlook
In future work, we would like to understand the relation between non-locality and fast scrambling better. This requires a careful investigation of scrambling time which is logarithmic in system size for fast scramblers, therefore requires investigation on system with large sizes. [^4] We leave this for future study due to the lack of resources to perform computation for higher dimensions. It would be interesting find the degree of non-locality require for fast scramblers by studying model with power-law decaying interactions. In ref PhysRevA.102.022402 , the authors shown that the fast scrambling is prohibited in models with a generic all-to-all term with prefactor $\sim 1/N^{\gamma}$ if $\gamma>1/2$. It would be interesting if this bound can be verified from operator growth using K-complexity.
Although all-to-all interactions are thought to be essential for fast scrambling belyansky2020minimal , certain systems demonstrate fast scrambling even in the absence of such interactions ikedaCommentsMinimalModel2021 . These models consider spin chains with next-to-nearest neighbor interactions in a specified combination of some sectors. The spread of quantum information in similar models with mild non-locality are studied in literature ecclesSpeedingSpreadQuantum2021a . These study show that although non-locality increases the rate of information spreading, but in lattice models these rates are suppressed. An extension of this work may consider K-complexity in such mild non-local models.
@@ Acknowledgements.
We wish to thank Sumilan Banerjee, Rathindra Nath Das, Bidyut Dey, Johanna Erdmenger, Mario Flory, Aaron Friedman, Michal Heller, Chethan Krishnan, Manas Kulkarni, Subroto Mukerjee, Pratiik Nandy, and Aninda Sinha for various useful discussions and comments about this and related works. The work of AB is supported by the Polish National Science Centre (NCN) grant 2021/42/E/ST2/00234.
@@ A.1 Level Statistics
In quantum chaos, Random Matrix Theory (RMT) is a powerful tool that accurately describes the spectral statistics of quantum systems whose classical counterparts exhibit chaotic behavior. For quantum Hamiltonians associated with integrable classical systems, the Berry-Tabor conjecture suggests that their level statistics follow a Poisson distribution 10.1007/978-3-0348-8266-8_36 .
However, for quantum Hamiltonians whose classical counterparts are chaotic, the Bohigas-Giannoni-Schmit conjecture proposes that their level statistics should fall into one of the three classical ensembles of RMT PhysRevLett.52.1 . These ensembles correspond to Hermitian random matrices with independently distributed entries as follows:
@@
The distribution $P(r)$ for random matrix ensembles Wigner ensembles (GOE, GUE, and GSE) is shown to PhysRevLett.110.084101
@@
Figure 9: Probability distribution functions of the $\tilde{r}$ statistics for fixed $Z$-reflection symmetry blocks and Parity block of the Hamiltonian (1), computed for $N=13$ spins in the $P=+1$, and $z=+1$ sector in (a) Integrable and (b) Chaotic limits. The results are obtained from averaging over 100 ensembles with $\epsilon$ drawn from normal distribution with mean 00 and deviation $10^{-4}$. The mean level spacing ratio obtained for the integrable and chaotic limits are 0.38733 and 0.53433 respectively.
Results
In Fig. 9, we showed the probability distribution functions of the $\tilde{r}$ statistics of the local TFIM Hamiltonian (1) in integrable and chaotic regimes. We compare the results for both $P(\tilde{r})$ and $\langle\tilde{r}\rangle$ with the analytical results for Poisson and GOE symmetry class. The numerical result are in good agreement with those of analytical.
@@ Results
In Fig. 10, we showed the probability distribution function of the $\tilde{r}$ statistics for non-local TFIM Hamiltonian (see Sec. 3.1) computed for $L=13$ spins. As reported in belyansky2020minimal , we find that non-locality term changes the integrable model to chaotic for $|\gamma|\gtrsim 0.25$.
Figure 11: Probability distribution functions of the $\tilde{r}$ statistics for Transverse mixed field Ising model for varying exponent $\alpha$. Refer Sec. 3.2 for parameter values.
In Fig. 11, we shown level statistics distribution for Transverse mixed field Ising model discussed in Sec 3.2 for varying non-local exponent $\alpha$. It is clear from here that it is hard to distinguish these different degrees of non-locality from the level statistics, while the initial growth rate of K-complexity can quite clearly distinguish them. The reason is that in K-complexity, we can distinguish between different degrees of non-locality during the pre-scrambling regime. At the same time, the level statistics can capture only the details about the nature of the spectrum, and there is no way to make such a timewise distinction.
@@ A.2 Spectral Form Factor
of the usual partition function $Z(\beta)=\text{Tr}\left(e^{-\beta H}\right)$. The function $Z(\beta,t)$ is an erratic function of time and, therefore oscillates largely at late times. The fluctuations are studied by considering the normalized squared quantity, the so-called spectral form factor (SFF) cotlerBlackHolesRandom2017 ; PhysRevD.95.126008 ; winerSpectralFormFactor2022
$g(\beta,t)=\left|\frac{Z(\beta,t)}{Z(\beta)}\right|^{2}=\frac{1}{Z(\beta)^{2}}\sum_{m,n}e^{-\beta(E_{m}+E_{n})}e^{i(E_{m}-E_{n})t}$ (26)
@@
$|\mathbb{E}\text{Tr}(e^{-iH(\beta+it)})|^{2}+\left(\mathbb{E}\left[|\text{Tr}(e^{-iH(\beta+it)})|^{2}\right]-|\mathbb{E}\text{Tr}(e^{-iH(\beta+it)})|^{2}\right)$ (30)
The first term is a disconnected part of the SFF, which comes solely from the average density of states. The second part is a connected part which contains the information on the correlation between energy levels. According to random matrix universality, an ensemble of quantum chaotic Hamilontians will generically have the same connected SFF as the canonical Gaussian ensembles of RMT. The conjectured universal profile of the SFF of the GUE contains three distinct regimes winerSpectralFormFactor2022 ; Cipolloni_2023 ; brezinSpectralFormFactor1997 :
@@
The overall profile for GOE and GSE symmetry class also follows a slope-dip-ramp-plateau pattern apart from a few details, such as the shape of the ramp and the plateau. For example, the SFF of the GUE symmetry class has a sharp kink, while in the GOE symmetry class, the kink is smoothed Cipolloni_2023 .

### 11 Mar 2024 v2

@@ III.3   Mixed field XXZ Hamiltonian
The XXZ Hamiltonian commutes with total spin operator $M$ in the $z$-direction and is also invariant under reflection with respect to the edge of the chain i.e. under parity operator $P$. In previous studies [49, 50, 51, 52, 53, 54, 55, 8], it was shown that the integrability can be broken by addition local term
@@
$\begin{split}\mathcal{H}^{(\alpha)}_{XXZ}&=\frac{1}{4}\sum_{i=1}^{N-1}J^{\alpha}_{ij}\,(\sigma_{i}^{x}\sigma_{i+1}^{x}+\sigma_{i}^{y}\sigma_{i+1}^{y})\\ &\qquad\qquad+\frac{\tilde{J}^{\alpha}_{ij}}{4}\,\sigma_{i}^{z}\sigma_{i+1}^{z}\end{split}$ (18)
@@
We study the initial Lanczos sequence for $L=12$ with varying exponent $\alpha$ for zero and non-zero value of $\epsilon_{d}$ (See Fig. 10 and 11). The initial operator $\mathcal{O}$ is chosen such that it respects the two symmetries of the model, i.e., parity and total spin in $z$-direction
@@ IV   Conclusions
IV.1   Discussion of results
Lanczos sequence The saturation and initial growth of Lanczos coefficients is greater in non-local TFIM Hamiltonian relative to the local case, which reflects the faster information scrambling in non-local systems. The local chaotic dynamics, along with long-range non-local interactions, are sufficient to give rise to fast scrambling. This is because the Lanczos sequence for chaotic Hamiltonian with local interactions saturates at a comparatively smaller value than its non-local chaotic counterpart. The large saturation value of Lanczos coefficients are the result of non-local interactions in the system, but they can equivalently be said to be the result of fast scrambling as well. The fluctuation in the non-local cases is also suppressed, which is a result of integrability breaking due to non-local interaction.This results in a higher saturation value of Krylov complexity.
Krylov Complexity Apart from common features of Krylov complexity, such as the transition from exponential growth to linear and the late time saturation, we find that the initial time growth rate is greater in non-local cases, indicating the onset of fast scrambling in the system. The higher growth rate in Krylov complexity for the non-local cases is a result of the initial faster growth of Lanczos coefficients in non-local cases due to the non-local interaction. Furthermore, since non-locality results in integrability breaking, we also find a large saturation value in Krylov complexity satisfying the bound in terms of the Krylov dimension. It is important to note that although the non-local interactions with the integrable local part of Hamiltonian show level statistics that of a chaotic model, it does not attain K-complexity saturation value similar to that of a chaotic case. In terms of associated Lanczos coefficients, the non-local integrable case has larger fluctuations as compared to the local chaotic case. It is noteworthy that even the introduction of the non-locality does not make the integrable model fully chaotic. On the other hand, when similar non-local terms are introduced in the otherwise local chaotic model, the late-time saturation value does not change much compared to the local chaotic model. However, the initial growth of complexity is faster in this case also. We reach similar conclusions through our study in the mixed field Ising and XXZ models. While the initial complexity growth acts as a probe of the degree of non-locality (different $\alpha$), the late-time saturation value only probes whether the underlying local theory ($\alpha>2$) is integrable or not.
@@ IV.1   Discussion of results
IV.2   Outlook
In future work, we would like to understand the relation between non-locality and fast scrambling better. This requires a careful investigation of scrambling time which is logarithmic in system size for fast scramblers, therefore requires investigation on system with large sizes. [^4] We leave this for future study due to the lack of resources to perform computation for higher dimensions. It would be interesting find the degree of non-locality require for fast scramblers by studying model with power-law decaying interactions. In ref [56], the authors shown that the fast scrambling is prohibited in models with a generic all-to-all term with prefactor $\sim 1/N^{\gamma}$ if $\gamma>1/2$. It would be interesting if this bound can be verified from operator growth using K-complexity.
Although all-to-all interactions are thought to be essential for fast scrambling [31], certain systems demonstrate fast scrambling even in the absence of such interactions [57]. These models consider spin chains with next-to-nearest neighbor interactions in a specified combination of some sectors. The spread of quantum information in similar models with mild non-locality are studied in literature [58]. These study show that although non-locality increases the rate of information spreading, but in lattice models these rates are suppressed. An extension of this work may consider K-complexity in such mild non-local models.
@@ Acknowledgements.
We wish to thank Sumilan Banerjee, Rathindra Nath Das, Bidyut Dey, Johanna Erdmenger, Mario Flory, Aaron Friedman, Michal Heller, Chethan Krishnan, Manas Kulkarni, Subroto Mukherjee, Pratik Nandy, and Aninda Sinha for various useful discussions and comments about this and related works. The work of AB is supported by the Polish National Science Centre (NCN) grant 2021/42/E/ST2/00234.
@@ A.1   Level Statistics
In quantum chaos, Random Matrix Theory (RMT) is a powerful tool that accurately describes the spectral statistics of quantum systems whose classical counterparts exhibit chaotic behavior. For quantum Hamiltonians associated with integrable classical systems, the Berry-Tabor conjecture suggests that their level statistics follow a Poisson distribution [35]. However, for quantum Hamiltonians whose classical counterparts are chaotic, the Bohigas-Giannoni-Schmit conjecture proposes that their level statistics should fall into one of the three classical ensembles of RMT [36]. These ensembles correspond to Hermitian random matrices with independently distributed entries as follows:
@@
The distribution $P(r)$ for random matrix ensembles Wigner ensembles (GOE, GUE, and GSE) is shown to [37]
@@
Figure 12: Probability distribution functions of the $\tilde{r}$ statistics for fixed $Z$-reflection symmetry blocks and Parity block of the Hamiltonian (1), computed for $L=13$ spins in the $P=+1$, and $z=+1$ sector in (a) Integrable and (b) Chaotic limits. The results are obtained from averaging over 100 ensembles with $\epsilon$ drawn from normal distribution with mean 00 and deviation $10^{-4}$. The mean level spacing ratio obtained for the integrable and chaotic limits are 0.38733 and 0.53433 respectively.
Results In Fig. 12, we showed the probability distribution functions of the $\tilde{r}$ statistics of the local TFIM Hamiltonian (1) in integrable and chaotic regimes. We compare the results for both $P(\tilde{r})$ and $\langle\tilde{r}\rangle$ with the analytical results for Poisson and GOE symmetry class. The numerical result are in good agreement with those of analytical.
@@
In Fig. 13, we showed the probability distribution function of the $\tilde{r}$ statistics for non-local TFIM Hamiltonian (see Sec. III.1) computed for $L=13$ spins. As reported in [31], we find that non-locality term changes the integrable model to chaotic for $|\gamma|\gtrsim 0.25$.
Figure 14: Probability distribution functions of the $\tilde{r}$ statistics for Transverse mixed field Ising model for varying exponent $\alpha$. Refer Sec. III.2 for parameter values.
In Fig. 14, we shown level statistics distribution for Transverse mixed field Ising model discussed in Sec III.2 for varying non-local exponent $\alpha$. It is clear from here that it is hard to distinguish these different degrees of non-locality from the level statistics, while the initial growth rate of K-complexity can quite clearly distinguish them. The reason is that in K-complexity, we can distinguish between different degrees of non-locality during the pre-scrambling regime. At the same time, the level statistics can capture only the details about the nature of the spectrum, and there is no way to make such a timewise distinction.
@@ A.2   Spectral Form Factor
of the usual partition function $Z(\beta)=\text{Tr}\left(e^{-\beta H}\right)$. The function $Z(\beta,t)$ is an erratic function of time and, therefore oscillates largely at late times. The fluctuations are studied by considering the normalized squared quantity, the so-called spectral form factor (SFF) [38, 39, 40]
$\begin{split}g(\beta,t)&=\left|\frac{Z(\beta,t)}{Z(\beta)}\right|^{2}\\ &=\frac{1}{Z(\beta)^{2}}\sum_{m,n}e^{-\beta(E_{m}+E_{n})}e^{i(E_{m}-E_{n})t}\end{split}$ (26)
@@
$\begin{split}&|\mathbb{E}\text{Tr}(e^{-iH(\beta+it)})|^{2}\\ &+\left(\mathbb{E}\left[|\text{Tr}(e^{-iH(\beta+it)})|^{2}\right]-|\mathbb{E}\text{Tr}(e^{-iH(\beta+it)})|^{2}\right)\end{split}$ (30)
The first term is a disconnected part of the SFF, which comes solely from the average density of states. The second part is a connected part which contains the information on the correlation between energy levels. According to random matrix universality, an ensemble of quantum chaotic Hamilontians will generically have the same connected SFF as the canonical Gaussian ensembles of RMT. The conjectured universal profile of the SFF of the GUE contains three distinct regimes [40, 59, 60]:
@@
The overall profile for GOE and GSE symmetry class also follows a slope-dip-ramp-plateau pattern apart from a few details, such as the shape of the ramp and the plateau. For example, the SFF of the GUE symmetry class has a sharp kink, while in the GOE symmetry class, the kink is smoothed [59].